A purse contains a collection of quaters, dimes,
nickels,and pennies.Thwe average value of the coins
in the purse is 18cents.If a penny is removed from the purse
the average value is___19.find how many nickels are there?
: A purse contains a collection of quaters, dimes,
nickels,and pennies.Thwe average value of the coins
in the purse is 18cents.If a penny is removed from the purse
the average value is___19.find how many nickels are there?

Answer by ankor@dixie-net.com(4683)

A purse contains a collection of quarters, dimes, nickels,and pennies.
The average value of the coins in the purse is 18 cents. If a penny is removed
from the purse, the average value is .19. Find how many nickels are there?
:
Let x = total value of the coins
Let y = no. of coins
:
x/y = .18
Or
x = .18y
then, when penny removed
((x-.01))/((y-1)) = .19
Substitute .18y for x:
((.18y-.01))/((y-1)) = .19
.18y - .01 = .19(y-1)
:
.18y - .01 = .19y - .19
:
+.19 - .01 = .19y - .18y
:
.18 = .01y
y = .18/.01
y = 18 coins
Find x
x = .18(18)
x = $3.24 total value
:
Check this: 3.23/17 = .19 (one penny removed)
:
Actually I don't think there will be any nickels.
:
In order to get this amt with 18 coins we must have:
12 quart = 3.00
2 dimes = 0.20
4 pennies =.04
----------------
18 coins = $3.24